In a beauty contest, the ages of the contestants are normally distributed with a mean of



24 years and a standard deviation of 1.5. What percent of the contestants have age



between 22.5 and 27 years old?

1
Expert's answer
2022-04-06T17:30:42-0400

"P(22.5<x<27)=P(\frac{22.5-24}{1.5}<z<\frac{27-24}{1.5})\\
=P(-1<z<2)\\
=P(0<z<1)+P(0<z<2)\\
=0.3413+0.4772=0.8185\\
the\; percentage\; is\;81.85\%"